40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 30 06:25:20 2011, kiwamu, Update, LSC, Calbiration of BS, ITMs and PRM actuators actuators.png
    Reply  Sat Oct 8 00:44:42 2011, kiwamu, Update, LSC, calibration of SRM actuator actuators.png
Message ID: 5583     Entry time: Fri Sep 30 06:25:20 2011     Reply to this: 5637
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: Calbiration of BS, ITMs and PRM actuators 
The AC responses of the BS, ITMs and PRM actuators have been calibrated.
 
(Background)
 To perform some interferometric works such as #5582, the actuator responses must be measured.
 
(Results)
     BS = 2.190e-08 / f2     [m/counts]
     ITMX  = 4.913e-09 / f2   [m/counts]
     ITMY  = 4.832e-09 / f2   [m/counts]
     PRM   = 2.022e-08 / f2  [m/counts]
 
actuators.png
 
(Measurement)
The same technique as I reported some times ago (#4721) were used for measuring the BS and ITMs actuators.
In order to measure the PRM actuator, power-recycled ITMY (PR-ITMY) was locked and the same measurement was applied.
The sensor response of PR-ITMY was calibrated by exciting the ITMY actuator since the response of the ITMY had been already measured.
ELOG V3.1.3-